OpenStudy (anonymous):

The sum of two numbers is 51. Their difference is 13. What is the smaller of the two numbers?

jhonyy9 (jhonyy9):

a+b=51 a-b=13 ------- 2a=64 a=32 32+b=51 b=51-32=19 a=32 b=19